OpenStudy (anonymous):

PLEASE HELP PLEASE IM ALMOST DONE One of Madison's hobbies is to locate hidden geocaches. To find yesterday's cache, Madison was instructed to start at a water fountain at a local park. Then, the directions had her walk 15 yards due west and 8 yards due south. If this was the location of the cache, how far was it from the water fountain where she started?

OpenStudy (anonymous):

A 23 yards B 17 yards C 12 yards

jigglypuff314 (jigglypuff314):

pythagorean theorem = √(15^2+8^2) = ?

OpenStudy (anonymous):

okay so how would I solve that

jigglypuff314 (jigglypuff314):

already told ya how... pythagorean theorem = √(15^2+8^2) = ?

OpenStudy (anonymous):

76

jigglypuff314 (jigglypuff314):

erm... try again?

OpenStudy (anonymous):

I don't have a scientific calculator

OpenStudy (anonymous):

could you please just answer this one for me plz

jigglypuff314 (jigglypuff314):

http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. so no, I will not give answers. 15*15 = ? 8*8 = ? then add them together then sqrt it

OpenStudy (anonymous):

8*8 is 49

OpenStudy (anonymous):

64

OpenStudy (anonymous):

u there

jigglypuff314 (jigglypuff314):

yes... 15*15=225 8*8 = 64 225 + 64 = 289 then sqrt 289 = ?

OpenStudy (anonymous):

how do I squre it

OpenStudy (anonymous):

@sajjadahmed

jigglypuff314 (jigglypuff314):

square root it

OpenStudy (anonymous):

its 17

OpenStudy (anonymous):

looks like you already got answer...